Giving a test to a group of students, the grades and gender are summarized below

A B C Total
Male 10 14 18 42
Female 19 11 15 45
Total 29 25 33 87



If one student is chosen at random,

Find the probability that the student was female AND got a "A".

Question 8 . 1 pts

A certain disease has an incidence rate of 0.8%. If the false negative rate is 8% and the false positive rate is 5%, compute the probability that a person who tests positive actually has the disease. round three decimal places.
Answer #1

1)

P( student was female AND got a "A".) =19/87

2)

P(tested positive)=P(have disease and tested positive)+P(not have disease and tested positive)
=0.008*0.92+0.992*0.05=0.057
Therefore P(have disease|tested positive)=0.008*0.92/0.057=0.129